a non uniform wire of length l and mass m has a variable linear mass density givenby u=kx where k is dist from one end of wire and k is constant . find the time taken by a pulse starting at one end to reach the other end when the tension in wire is T.


 

Grade:

1 Answers

Abhishek Kumar
9 years ago
v = \sqrt{T/kx} at x distance.
Time taken in travelling the dx distance at x.
dt = dx/v
So, T = Integral (\sqrt{kx/T} dx)